OpenStudy (anonymous):

Solve for x. 3:12 = x:16 A. 2 1/4 B. 4 C. 64

mathslover (mathslover):

\[\huge{\frac{3}{12}=\frac{x}{16}}\] You can easily solve this by cross multiplying : \[\huge{3*16=12*x}\]

mathslover (mathslover):

Can you solve this now ? for x

Parth (parthkohli):

\(\mathbf{\LaTeX}\) tip ----> \(\times\) or \(\cdot\) can be used for multiplication.

OpenStudy (anonymous):

its B.

mathslover (mathslover):